OpenStudy (anonymous):

∑[i=1,∞,(12)(0.25)^i]

OpenStudy (anonymous):

\[12\sum_{i=1}^{\infty}\left(\frac{1}{4}\right)^i\] is a start

OpenStudy (anonymous):

you are starting at \(i=1\) not \(i=0\) right?

OpenStudy (anonymous):

correct. & i am to find the sum of this infinite geometric series.

OpenStudy (anonymous):

correct. & i am to find the sum of this infinite geometric series.

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty}r^i=\frac{r}{1-r}\] so \[\sum_{i=1}^{\infty}(\frac{1}{4})i=\frac{\frac{1}{4}}{1-\frac{1}{4}}\] \[=\frac{\frac{1}{4}}{\frac{3}{4}}=\frac{1}{3}\] and \[12\times\frac{1}{3}=4\]

OpenStudy (anonymous):

Thank you SO much.

OpenStudy (anonymous):

yw